# Data Assimilation with Machine Learning Surrogate Models: A Case Study with FourCastNet

### v1

@@ Abstract
Modern data-driven surrogate models for weather forecasting provide accurate short-term predictions but inaccurate and nonphysical long-term forecasts. This paper investigates online weather prediction using machine learning surrogates supplemented with partial and noisy observations. We empirically demonstrate and theoretically justify that, despite the long-time instability of the surrogates and the sparsity of the observations, filtering estimates can remain accurate in the long-time horizon. As a case study, we integrate FourCastNet, a state-of-the-art weather surrogate model, within a variational data assimilation framework using partial, noisy ERA5 data. Our results show that filtering estimates remain accurate over a year-long assimilation window and provide effective initial conditions for forecasting tasks, including extreme event prediction.
@@ 1 Introduction
The potential of weather surrogate models to accelerate extreme event prediction has recently received attention at the government level, with an April 2024 U.S. Executive Office report calling attention to its potential widespread operational use [The President’s Council of Advisors on Science and Technology, 2024]. This report stated that in the U.S. in 2023 alone, the economic damage due to extreme weather events totalled to \$92.9 billion from 28 weather disasters, and the frequency of these extreme events is expected to continue to increase in the coming years [The President’s Council of Advisors on Science and Technology, 2024]. This staggering economic loss and the projected increased frequency in weather disasters point to the escalating need for real-time, accurate weather forecasting. Surrogate weather models allow for real-time forecasting and uncertainty quantification given an accurate estimate of the initial or current weather state, which could be especially impactful for time-sensitive extreme events like hurricanes, typhoons, cyclones, etc. In the event of extreme weather, it is paramount to produce, in real time, accurate estimates of the state of the weather to initialize forecasts, which can be accomplished via data assimilation. Moreover, quickly computing large ensembles of forecasts using these weather surrogates allows for a timely characterization of the distribution of potential future outcomes. Speed, accuracy, and uncertainty quantification are crucial to gauge the severity and likelihood of these potential outcomes and subsequently inform timely decisions regarding public safety. In our evaluation tasks for extreme events, we especially inspect the predicted severity of these events compared to the actual severity. As a case study, we compute forecasts of Typhoon Mawar in 2023 and assess the U- and V-component wind speed at 10 meters (10m) above Earth’s surface, the mean sea level pressure, and the location of the typhoon’s eye to characterize the quality of forecasts given various types of initializations, including the data assimilation analyses we produce.
@@
In our work, we aim to assess the utility of weather surrogate models, particularly a state-of-the-art weather surrogate model, FourCastNet [Pathak et al., 2022], in the three main tasks of operational interest: (1) estimation of the high-dimensional weather state from low-resolution observations via a variational data assimilation framework, (2) multi-step-ahead forecasting using these estimates, and as a more focused forecasting task, (3) extreme event prediction using these estimates. Our high-dimensional state consists of 20 atmospheric features at various pressure levels in the atmosphere at a global $0.25^{\circ}$ resolution, corresponding to a $720\times 1440$ grid. We utilize (1) FourCastNet as our weather model, (2) low-resolution, noisy ERA5 reanalysis data [Hersbach et al., 2020] as a proxy for real observations, and (3) a 3DVar variational data assimilation algorithm [Lorenc et al., 2000] to combine forecasts and observations. We demonstrate that pairing FourCastNet within 3DVar to assimilate low-resolution, noisy observational data can produce visually realistic weather patterns while maintaining a stable reconstruction error over a long time horizon, and we provide a theoretical justification for this result. Additionally, we show that these 3DVar analyses can serve as effective initial conditions for forecasting tasks, even in our extreme event case study.
@@ 2 Contributions
This work provides new empirical and theoretical evidence that data-driven surrogate models for weather forecasting can be successful in data assimilation tasks. We illustrate the potential of purely data-driven, global weather data assimilation using a current state-of-the-art weather surrogate model, FourCastNet, and low-resolution, noisy data. Our assimilation is based on a simple 3DVar filter that can be run with a single 80GB NVIDIA A100 GPU. The main contributions of this paper are to:
@@ 3 Related work
In recent years, there have been substantial advances in machine learning to create surrogates of numerical weather models in order to produce predictions at a fraction of the cost while maintaining short-term accuracy. Notable global data-driven weather models include FourCastNet [Pathak et al., 2022], Pangu [Bi et al., 2023], GraphCast [Lam et al., 2023], a graph-based weather model from [Keisler, 2022], FengWu [Chen et al., 2023a], and FuXi [Chen et al., 2023b]. In this work, we focus on evaluating FourCastNet as a case study.
@@
The successes on smaller scale data point to potential success for larger scale integration of weather surrogates in data assimilation tasks. Huang et al. [2024] and Xiao et al. [2023] are among the first works to combine data assimilation and a state-of-the-art weather surrogate at a high dimensional scale. Huang et al. [2024] performs a data assimilation task at an operational scale by reconstructing 24 atmospheric features at a $0.25^{\circ}$ resolution (721$\times$1440 grid) with the data-driven weather surrogate GraphCast [Lam et al., 2023]. In addition, the authors formulate a novel alternative to traditional data assimilation tasks through the use of a learnable diffusion model. Huang et al. [2024], however, focuses on short assimilation horizons and one-step-ahead forecasting in its experiments and utilizes a diffusion-based architecture, which may suffer from inefficiencies in sample generation speed. Addressing this generation speed issue is an area of active research [Chen et al., 2024]. In our work, we expand upon evaluation tasks to assess the quality of longer forecasts using data assimilation analyses in a computationally efficient manner.
@@
Xiao et al. [2023] combines FenguWu [Chen et al., 2023a] within a 4DVar assimilation scheme to estimate a latent representation consisting of 69 atmospheric features at a roughly $1.41^{\circ}$ spatial resolution, corresponding to a $128\times 256$ grid. This work assimilates observations that were created from ERA5 data with randomly applied masks, leading to 15% of locations being observed [Xiao et al., 2023]. Similarly to the results we present in this paper, Xiao et al. [2023] shows stable analysis errors across the span of a year. Our work demonstrates similar error stability for a higher dimensional representation of global weather patterns (0.25 resolution) using lower resolution observations (observing at most 1.5% of locations) compared to Xiao et al. [2023] for observations acquired at fixed spatial locations.
@@ 4 Data description
ERA5 is a reanalysis dataset that provides hourly atmosphere, land, and ocean feature estimates produced by the variational data assimilation method 4DVar [Rawlins et al., 2007] at a resolution of 0.25 using observational weather data from 1979 to present day [Hersbach et al., 2020]. Data was pulled from the Copernicus Climate Data Store, Sabater [2019] for land features, and Hersbach et al. [2023] for pressure level features. We retain only a subset of the atmospheric features in ERA5, specifically total column water vapor (TCWV), geopotential at 50, 500, 850, and 1000 hPa, U-component wind speed at 10 meters from the surface and at 500, 850, and 1000 hPa, V-component wind speed at 10 meters from the surface and at 500, 850, and 1000 hPa, relative humidity at 500 and 850 hPa, temperature at 2 meters from the surface and at 500 and 850 hPa, surface pressure (sp), and mean sea level pressure (mslp). Consequently, we retained all the features that FourCastNet was trained to predict. Additionally, we standardized the ERA5 dataset using the same global feature means and standard deviations that were used in training FourCastNet.
As described next, for our assimilation and forecasting tasks, the true latent states represent high-resolution ERA5 data across 2023, from which we create low-resolution noisy observations. We emphasize that our chosen time range, ERA5 data for 2023, is disjoint from the time range that FourCastNet was trained and tuned on, ERA5 data from 1979 to 2017 [Pathak et al., 2022].
True latent states $\{x_{t}^{\text{\rm true}}\}_{t=0}^{T}$.
The true state $x_{t}^{\text{true}}\in\mathbb{R}^{d_{x}}$ at time $t$ represents 20 ERA5 atmospheric features at a $0.25^{\circ}$ resolution, which results in a state dimension $d_{x}=20\times 720\times 1440=20,736,000$ at one time point. This state is partially observed every 6 hours for the entirety of 2023, so we consider $T=1460$ time points (4 observations/day× $\ \times\ 365\text{ days}$). Thus, $t=0$ corresponds to January 1, 2023 at 00:00 UTC, and $t=T$ corresponds to December 31, 2023 at 18:00 UTC.
@@
Table 1: Table describing the observational dataset resolution with corresponding distances between observations along the equator and percentage of the latent states observed for each dataset. In each of these datasets, we observe all 20 atmospheric features for every $k$-th latent location in both the latitude and longitude directions. As reference, the latent states have a resolution of 0.25, which corresponds to a distance of 28 km between latent states along the equator and a latitude/longitude grid size of 720$\times$1440.
@@ 5.1 Setting
Our goal is to estimate a high-dimensional latent state $\{x_{t}^{\text{true}}\}_{t\geq 1}$ given observations $\{y_{t}\}_{t\geq 0}$, which are derived from the following setting:
@@
The observation operator $H\in\{0,1\}^{d_{y}\times d_{x}}$ is a subset of the rows of the identity matrix, with the remaining rows indicating wherever a latent dimension is observed. Our $H$ is independent of time, meaning that we observe the same subset of locations throughout the entire assimilation horizon.
@@
To enhance filter stability, we utilize a smoothing operator $S$ defined as a 2D convolution with weight matrix $W^{(4)}\in\mathbb{R}^{4\times 4}$ as in equation (5), and stride $(1,1)$ for each of the 20 atmospheric features. We set $\sigma^{2}=8$. This smoothing operation attempts to control instabilities in the dynamics model.
@@
We specify that $C=qBB^{\text{T}}$, where $B$ is a matrix representing 2D convolution with weight matrix $W^{(k)}\in\mathbb{R}^{k\times k}$ defined in equation (5), and stride $(1,1)$ for each of the 20 atmospheric features. In our experiments, we vary the size of the convolutional kernel across observational data resolutions according to $k$. In our experiments, we choose $q=0.5/\sum_{i=1}^{k}\sum_{j=1}^{k}\{W^{(k)}_{i,j}\}^{2}$. The constant 0.5 was heuristically chosen to be a similar magnitude to one-step-ahead forecasting errors for $\mathcal{F}_{s}$ in the standardized space. For each $W^{(k)}$, we set $\sigma^{2}=8$. With our choices of $C$ and $H$ for each observation resolution, the matrix $(HCH^{\text{T}}+R)^{-1}$ in the analysis step in (3) is diagonal, which results in substantial computational savings. This $C$ matrix was constructed mainly to maximize computational efficiency and may lead to some physically unrealistic analyses that cause FourCastNet predictions to degrade. Future work can include a more sophisticated construction of this $C$, for example, via the widely adopted National Meteorological Center’s method described in Parrish and Derber [1992].
@@ 5.3 Theoretical long-time accuracy of 3DVar
We include a proof of Theorem 1 in Appendix B.
@@ 6 Results
Our results evaluate the empirical long-term assimilation stability of 3DVar with our chosen surrogate $\mathcal{F}_{s}$ in (6) and with varying resolutions of noisy ERA5 data as observations. We evaluate the forecasting performance using our 3DVar analyses as initial conditions and compare against a more naive approach of forecasting using only the interpolated observations as initial conditions. These interpolated observations are constructed in the same way as our 3DVar initialization $x_{0}$ in Section 5.2 for each $\{y_{t}\}_{t=1}^{T}$. The performance of these two approaches is averaged across 20 standardized atmospheric features and compared to a so-called “gold standard” where we compute forecasting metrics using ground truth ERA5 initializations. The comparison to this gold standard is included to contextualize how well we could expect to perform in these forecasting tasks in an ideal setting: ground truth information upon initialization. To further explore the task of forecasting, we assess the forecasting performance of an extreme event, Category 5-equivalent super typhoon, Typhoon Mawar in 2023.
We evaluate each of our tasks on 2023 ERA5 reanalysis data using the metrics latitude-weighted root-mean-square-error (RMSE) and latitude-weighted anomaly correlation coefficient (ACC). We provide a detailed explanation of our error metrics in Appendix C.

### v2

@@ Abstract
Modern data-driven surrogate models for weather forecasting provide accurate short-term predictions but inaccurate and nonphysical long-term forecasts. This paper investigates online weather prediction using machine learning surrogates supplemented with partial and noisy observations. We empirically demonstrate and theoretically justify that, despite the long-time instability of the surrogates and the sparsity of the observations, filtering estimates can remain accurate in the long-time horizon. As a case study, we integrate FourCastNet, a weather surrogate model, within a variational data assimilation framework using partial, noisy ERA5 data. Our results show that filtering estimates remain accurate over a year-long assimilation window and provide effective initial conditions for forecasting tasks, including extreme event prediction.
@@ 1 Introduction
The potential of weather surrogate models to accelerate extreme event prediction has recently received attention at the government level, with an April 2024 U.S. Executive Office report calling attention to its potential widespread operational use [The President’s Council of Advisors on Science and Technology, 2024]. This report stated that in the U.S. in 2023 alone, the economic damage due to extreme weather events totalled to \$92.9 billion from 28 weather disasters, and the frequency of these extreme events is expected to continue to increase in the coming years [The President’s Council of Advisors on Science and Technology, 2024]. This staggering economic loss and the projected increased frequency in weather disasters point to the escalating need for real-time, accurate weather forecasting.
Surrogate weather models allow for real-time forecasting and uncertainty quantification given an accurate estimate of the initial or current weather state, which could be especially impactful for time-sensitive extreme events like hurricanes, typhoons, cyclones, etc. In the event of extreme weather, it is paramount to produce, in real time, accurate estimates of the state of the weather to initialize forecasts, which can be accomplished via data assimilation. Moreover, quickly computing large ensembles of forecasts using these weather surrogates allows for a timely characterization of the distribution of potential future outcomes. Speed, accuracy, and uncertainty quantification are crucial to gauge the severity and likelihood of these potential outcomes and subsequently inform timely decisions regarding public safety.
@@
In our work, we aim to assess the utility of weather surrogate models, particularly a weather surrogate model, FourCastNet [Pathak et al., 2022], in the three main tasks of operational interest: (1) estimation of the high-dimensional weather state from low-resolution observations via a variational data assimilation framework, (2) multi-step-ahead forecasting using these estimates, and as a more focused forecasting task, (3) extreme event prediction using these estimates. Our high-dimensional state consists of 20 atmospheric features at various pressure levels in the atmosphere at a global $0.25^{\circ}$ resolution, corresponding to a $720\times 1440$ grid. We utilize (1) FourCastNet as our weather model, (2) low-resolution, noisy data derived from the ERA5 reanalysis [Hersbach et al., 2020] as a proxy for real observations, and (3) a 3DVar variational data assimilation algorithm [Lorenc et al., 2000] to combine forecasts and observations. We demonstrate that pairing FourCastNet within 3DVar to assimilate low-resolution, noisy observational data can produce visually realistic weather patterns while maintaining a stable reconstruction error over a long time horizon, and we provide a theoretical justification for this result. Additionally, we show that these 3DVar analyses can serve as effective initial conditions for forecasting tasks, even in our extreme event case study.
@@ 2 Contributions
This work provides new empirical and theoretical evidence that data-driven surrogate models for weather forecasting can be successful in data assimilation tasks. We illustrate the potential of purely data-driven, global weather data assimilation using a current weather surrogate model, FourCastNet, and low-resolution, noisy data. Our assimilation is based on a simple 3DVar filter that can be run with a single NVIDIA A100 GPU. The main contributions of this paper are to:
@@ 3 Related work
In recent years, there have been substantial advances in machine learning to create surrogates of numerical weather models in order to produce predictions at a fraction of the cost while maintaining short-term accuracy. Notable global data-driven weather models include FourCastNet versions 1 and 2 [Pathak et al., 2022; Bonev et al., 2023b], Pangu-Weather [Bi et al., 2023], GraphCast [Lam et al., 2023], a graph-based weather model from [Keisler, 2022], FengWu [Chen et al., 2023a], and FuXi [Chen et al., 2023b]. In this work, we focus on evaluating FourCastNet as a case study.
@@
The successes on smaller scale data point to potential success for larger scale integration of weather surrogates in data assimilation tasks. Huang et al. [2024] and Xiao et al. [2023] are among the first works to combine data assimilation and a weather surrogate at a high dimensional scale. Huang et al. [2024] performs a data assimilation task at an operational scale by reconstructing 24 atmospheric features at a $0.25^{\circ}$ resolution (721$\times$1440 grid) with the data-driven weather surrogate GraphCast [Lam et al., 2023]. In addition, the authors formulate a novel alternative to traditional data assimilation tasks through the use of a learnable diffusion model. Huang et al. [2024], however, focuses on short assimilation horizons and one-step-ahead forecasting in its experiments and utilizes a diffusion-based architecture, which may suffer from inefficiencies in sample generation speed. Addressing this generation speed issue is an area of active research [Chen et al., 2024]. In our work, we expand upon evaluation tasks to assess the quality of longer forecasts using data assimilation analyses in a computationally efficient manner.
@@
Xiao et al. [2023] combines FenguWu [Chen et al., 2023a] within a 4DVar assimilation scheme to estimate the state of the atmosphere for 69 atmospheric features at a roughly $1.41^{\circ}$ spatial resolution, corresponding to a $128\times 256$ grid. This work assimilates observations that were created from ERA5 data with randomly applied masks, leading to 15% of locations being observed [Xiao et al., 2023]. Similarly to the results we present in this paper, Xiao et al. [2023] shows stable analysis errors across the span of a year. Our work demonstrates similar error stability for a higher dimensional representation of global weather patterns (0.25 resolution) using lower resolution observations (observing at most 1.5% of locations) compared to Xiao et al. [2023] for observations acquired at fixed spatial locations.
Extreme event forecasting using data-driven weather surrogates.
As numerous cheap-to-evaluate machine learning weather surrogates have been developed and released to the public in recent years, preliminary research efforts have begun to assess the forecast quality in predicting extreme events, including hurricane, cyclone, and typhoon tracking. To highlight a few of these works and their main takeaways regarding extreme storm forecasting, Magnusson [2023] notes that Pangu-Weather’s forecasts of cyclone Eunice in 2022 were lacking small-scale features in the atmospheric fields, under-predicted the maximum wind speed of the cyclone, and showed a faster evolution of the cyclone compared to the NWP forecasts. Lopez-Gomez et al. [2023] notes that training their data-driven weather surrogate with a mean-squared-error loss, which is commonly used to train weather surrogate models, produced less skillful forecast of extremes and overly-smoothed forecasts, especially for larger lead times. Lastly, Charlton-Perez et al. [2024] notes that the weather surrogates FourCastNet, FourCastNet v2, Pangu-Weather, and GraphCast failed to produce small scale bands of strong winds in their Storm Ciarán case study and consistently under-predicted wind speed; however, the models showed comparable skill to NWP forecasts up to a 48 hour lead time.
Two key takeaways have emerged from this expanding body of literature, evidenced by the aforementioned work and others: machine learning weather surrogates generally produce overly smoothed predictions, and these models generally under-predict the intensity of extreme events. This issue is actively being investigated, and diffusion-based generative models like GenCast [Price et al., 2024] have shown promise in increasing sharpness and predicting closer to the true extremity of rare events, though at a substantial computational cost. In our evaluation of a particular extreme event, we especially inspect the predicted versus actual severity, as well as the level of detail in the predictions. As a case study, we compute forecasts of Typhoon Mawar in 2023 and assess the U- and V-component wind speed at 10 meters (10m) above Earth’s surface, the mean sea level pressure, and the location of the typhoon’s eye to characterize the quality of forecasts given various types of initializations, including the data assimilation analyses we produce.
@@ 4 Data description
4.1 ECMWF Reanalysis v5 (ERA5)
ERA5 is a reanalysis dataset that provides hourly atmosphere, land, and ocean feature estimates produced by the variational data assimilation method 4DVar [Rawlins et al., 2007] at a resolution of 0.25 using observational weather data from 1979 to present day [Hersbach et al., 2020]. Data was pulled from the Copernicus Climate Data Store, Sabater [2019] for land features, and Hersbach et al. [2023] for pressure level features. We retain only a subset of the atmospheric features in ERA5, specifically total column water vapor (TCWV), geopotential at 50, 500, 850, and 1000 hPa, U-component wind speed at 10m from the surface and at 500, 850, and 1000 hPa, V-component wind speed at 10m from the surface and at 500, 850, and 1000 hPa, relative humidity at 500 and 850 hPa, temperature at 2 meters from the surface and at 500 and 850 hPa, surface pressure (sp), and mean sea level pressure (mslp). Consequently, we retained all the features that FourCastNet was trained to predict. Additionally, we standardized the ERA5 dataset using the same global feature means and standard deviations that were used in training FourCastNet.
As described next, for our assimilation and forecasting tasks, the states we attempt to estimate are the high-resolution ERA5 data across 2023 in its native 0.25 resolution, from which we create low-resolution noisy observations. We emphasize that our chosen time range, ERA5 data for 2023, is disjoint from the time range that FourCastNet was trained and tuned on, ERA5 data from 1979 to 2017 [Pathak et al., 2022].
Ground truth states
$\{x_{t}^{\text{\rm true}}\}_{t=0}^{T}$. The ground truth state $x_{t}^{\text{true}}\in\mathbb{R}^{d_{x}}$ at time $t$ represents 20 atmospheric features at a $0.25^{\circ}$ resolution, which results in a state dimension $d_{x}=20\times 720\times 1440=20,736,000$ at one time point. This state is partially observed every 6 hours for the entirety of 2023, so we consider $T=1460$ time points (4 observations/day× $\ \times\ 365\text{ days}$). Thus, $t=0$ corresponds to January 1, 2023 at 00:00 UTC, and $t=T$ corresponds to December 31, 2023 at 18:00 UTC. Since we do not ever observe the true ground truth state in reality, we will evaluate our results based on a state-of-the-art proxy of atmospheric states, namely the ERA5 dataset [Hersbach et al., 2020].
@@ 4.1 ECMWF Reanalysis v5 (ERA5)
Table 1: Table describing the observational dataset resolution with corresponding distances between observations along the equator and percentage of the states observed for each dataset. In each of these datasets, we observe all 20 atmospheric features for every $k$-th location of interest in both the latitude and longitude directions. As reference, these states have a resolution of 0.25, which corresponds to a distance of 28 km between states along the equator and a latitude/longitude grid size of 720$\times$1440.
4.2 High-resolution forecasts (HRES) of the European Centre for Medium-Range Weather Forecasts (ECMWF)’s Integrated Forecasting System (IFS)
In Section 66.3, we utilize an additional dataset: archived high-resolution forecasts of ECMWF’s IFS, which we refer to as IFS-HRES. For our forecasting task of Typhoon Mawar, we pulled forecasts at 6 hour intervals of IFS-HRES initialized on May 23, 2023 at 00:00 UTC until May 30, 2023 at 12:00 UTC. The native resolution of the IFS-HRES forecasts are 0.1, but in order to match the resolution of the ERA5 data, we pulled these forecasts at 0.25. Since we use these forecasts in our analysis of Typhoon Mawar, we only retained atmospheric variables relevant to this prediction task, which include mean sea level pressure and U and V component wind speeds at 10m above the surface.
4.3 Observational typhoon data from the International Best Track Archive for Climate Stewardship (IBTrACS)
Section 66.3 additionally utilizes IBTrAC observational data from the Joint Typhoon Warning Center, which contains key information about tropical cyclones including their locations and intensities over time. We subset this data to focus specifically on Typhoon Mawar between May 23, 2023 00:00 UTC and May 30, 2023 12:00 UTC.
@@ 5.1 Setting
Our goal is to estimate a high-dimensional gridded representation of atmospheric features, denoted $\{x_{t}^{\text{true}}\}_{t\geq 1}$, given observations $\{y_{t}\}_{t\geq 0}$, which are derived from the following setting:
@@
The observation operator $H\in\{0,1\}^{d_{y}\times d_{x}}$ is a subset of the rows of the identity matrix, with the remaining rows indicating wherever a state coordinate is observed. Our $H$ is independent of time, meaning that we observe the same subset of locations throughout the entire assimilation horizon.
@@
A known limitation of FourCastNet is its forecasting instability near the poles [Bonev et al., 2023a]. To enhance filter stability, we utilize a smoothing operator $S$ defined as a 2D convolution with weight matrix $W^{(4)}\in\mathbb{R}^{4\times 4}$ as in equation (5), and stride $(1,1)$ for each of the 20 atmospheric features. We set $\sigma^{2}=8$. This smoothing operation attempts to control instabilities in the dynamics model. Appendix B provides an example visualization showcasing filter divergence in its early stage when smoothing is not applied to FourCastNet’s forecasts within 3DVar, assimilating 4.5 observations.
@@
We specify that $C=qBB^{\text{T}}$, where $B$ is a matrix representing 2D convolution with weight matrix $W^{(k)}\in\mathbb{R}^{k\times k}$ defined in equation (5), and stride $(1,1)$ for each of the 20 atmospheric features. In our experiments, we vary the size of the convolutional kernel across observational data resolutions according to $k$. In our experiments, we choose $q=0.5/\sum_{i=1}^{k}\sum_{j=1}^{k}\{W^{(k)}_{i,j}\}^{2}$. The constant 0.5 was heuristically chosen to be a similar magnitude to one-step-ahead forecasting errors for $\mathcal{F}_{s}$ in the standardized space. For each $W^{(k)}$, we set $\sigma^{2}=8$. With our choices of $C$ and $H$ for each observation resolution, the matrix $(HCH^{\text{T}}+R)$ in the analysis step in (3) is diagonal, which avoids $d_{y}^{3}$ operations for a matrix inversion and instead computes the inverse of $d_{y}$ scalars, resulting in substantial computational savings. For example when $k=8$, avoiding the matrix inversion reduces the number of computations from on the order of $10^{16}$ operations to on the order of $10^{5}$ operations.
This $C$ matrix was constructed mainly to maximize computational efficiency and may lead to some physically unrealistic analyses that cause FourCastNet predictions to degrade. Future work can include a more sophisticated construction of this $C$, for example, via the widely adopted National Meteorological Center’s method described in Parrish and Derber [1992].
@@ 5.3 Theoretical long-time accuracy of 3DVar
We include a proof of Theorem 1 in Appendix D.
@@ 6 Results
Our results evaluate the empirical long-term assimilation stability of 3DVar with our chosen surrogate $\mathcal{F}_{s}$ in (6) and with varying resolutions of noisy ERA5 data as observations. We evaluate the forecasting performance using our 3DVar analyses as initial conditions and compare against a more naive approach of forecasting using only the interpolated observations as initial conditions. These interpolated observations are constructed in the same way as our 3DVar initialization $x_{0}$ in Section 55.2 for each $\{y_{t}\}_{t=1}^{T}$. The performance of these two approaches is averaged across 20 standardized atmospheric features and compared to an idealized setting where we compute forecasting metrics using ground truth ERA5 initializations. We include this setting to contextualize how well we could expect to perform in these forecasting tasks in an ideal setting: ground truth information upon initialization. To further explore the task of forecasting, we assess the forecasting performance of an extreme event, Category 5-equivalent super typhoon, Typhoon Mawar in 2023.
We evaluate each of our tasks on 2023 ERA5 reanalysis data using the metrics latitude-weighted root-mean-square-error (RMSE) and latitude-weighted anomaly correlation coefficient (ACC) on ERA5’s native 0.25 resolution. We provide a detailed explanation of our error metrics in Appendix E.